Circular No. 7301

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999es IN ANONYMOUS GALAXY J. Mueller reports her discovery of a supernova (mag ∼ 18.5) on a red

plate taken on Nov. 2 UT by K. Rykoski and herself with the 1.2-m Oschin Schmidt Telescope in the course of the Palomar Outer Solar System Ecliptic Survey. SN 1999es is located at α = 2^{h}32^{m}43^{s}.4, δ = +22^{o}27′08′′ (equinox 2000.0), which is 12′′.8 west and 2′′ north of the host galaxy. There is no object at the position of SN 1999es on the Digitized Sky Survey or on a red plate from the second Palomar Sky Survey. A spectrum of the SN 1999es, obtained by M. Sawicki, H. Lin, T. Galema, and R. Burruss on Nov. 4 with the Palomar 5-m Hale Telescope (+ double spectrograph), looks like that of a type-Ia supernova, ∼ 2 weeks after maximum. The spectrum gives a tentative redshift of 0.07.

GRB 991105 M. J. S. Smith and J. in’t Zand, Space Research Organization Nether-

lands, Utrecht; G. Tarei and M. Dadina, BeppoSAX Scientific Operation Center, Rome; and L. Piro, Istituto di Astrofisica Spaziale, CNR, Rome, report: “The BeppoSAX Gamma Ray Burst Monitor (GRBM) and Wide Field Camera (WFC unit 2) observed a γ-ray burst triggered on Nov. 5.6950 UT (BATSE trigger 7701). A preliminary analysis reveals a duration of 13 s in the GRBM, with a peak intensity of 970 counts/s in the energy band 40–700 keV. In the WFC, the duration is 40 s and the peak flux is 1.1 Crab (2–28 keV). The position of the x-ray counterpart is α = 12^{h}03^{m}01^{s}, δ = −66^{o}48′49′′ (equinox 2000.0), with a 99-percent error radius of 3′.6. A BeppoSAX NFI follow-up observation will not be performed due to sun-angle constraints.”

SUPERNOVA 1999ek IN UGC 3329 L. G. Strolger, R. C. Smith, and N. B. Suntzeff, Cerro Tololo Interamer-

ican Observatory (CTIO); M. Hamuy, University of Arizona; and M. M. Phillips, Las Campanas Observatory, report: “Spectra (range 500–900 nm) obtained by Smith on Oct. 25 UT, and by P. Ugarte (CTIO) on Oct. 26, at the CTIO 1.5-m telescope indicate that SN 1999ek is a type-Ia supernova that was then at or near peak brightness. Photometric observations made by L. González (University of Chile) on Oct. 26, and by Hamuy and M. Wischnjewsky on Oct. 27, at the CTIO 0.9-m telescope show an increase in brightness of ∼ 0.22 mag in B, and ∼ 0.08 mag in V.”

1999 November 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green